Yes No Girls 42 ,110 Boys 95, 53 a. If b represents the event of selecting...

Yes No

Girls 42 ,110

Boys 95, 53

a. If b represents the event of selecting a response from a boy, find the complement of b [P(Ђ)]. 148/300=0.493 1-0.493=0.507 ?

b. Find the probability that the selected answer was no [P(no)]. ?

c. If g represents the event of selecting a response from a girl, find the complement of g [P (ḡ)]. 0.493?

d. Find the probability that the selected answer was yes [P(yes)].?
Answer #1

Part a is perfect

b. Find the probability that the selected answer was no

[P(no)] = count(total no) / count (total)

= (110+53)/300

= 163/300

= 0.5433

part c is perfect

d. Find the probability that the selected answer was yes [P(yes)].?

=> P(yes) = 1 - P(no)

= 1 - 0.5433

= 0.4567